# ModernVBERT: Towards Smaller Visual Document Retrievers

### v1

@@ Abstract
Multimodal embedding models are gaining prevalence, notably for document retrieval as efficient alternatives to text-only pipelines. These models are typically built by finetuning large vision–language decoders (VLMs) with contrastive losses on text–image pairs. In this work, we show that, while cost-efficient, this repurposing approach often bottlenecks retrieval performance. Through controlled experiments, we establish a principled recipe for improving visual document retrieval models. We notably measure the impact of attention masking, image resolution, modality alignment data regimes, and late interaction centered contrastive objectives which emerge as central performance factors. Building on these insights, we release ModernVBERT, a compact 250M-parameter vision–language encoder that outperforms models up to 10 times larger when finetuned on document retrieval tasks. Models and code are made available at https://huggingface.co/ModernVBERT.
@@ 1 Introduction
A core challenge in visual retrieval is learning text–image embeddings that faithfully encode complex, multi-level semantics—from fine-grained details to high-level concepts. For many years, contrastive pre-training of dual encoders, typically pairing a vision transformer (ViT) (Dosovitskiy et al., 2021) with a text encoder, was the standard approach for generating cross-modal embeddings (Radford et al., 2021; Zhai et al., 2023). Owing to the performance improvements enabled by generative vision–language models (VLMs), recent work has increasingly explored repurposing these models as multimodal encoders through extended contrastive post-training (Ma et al., 2024; Faysse et al., 2025; Jiang et al., 2025).
While this approach capitalizes on the inherent capabilities of the backbone generative model to address complex visual tasks, it is unclear whether the design decisions made for generative purposes hinder the models when repurposed for representation learning. Typically, VLM-based encoders inherit the causal attention masks VLMs are trained with, which has been repeatedly shown to be suboptimal for embedding tasks (Li et al., 2023; BehnamGhader et al., 2024; Lee et al., 2025; Springer et al., 2025; Gisserot-Boukhlef et al., 2025). Another notable architectural property of contemporary VLMs is their high parameter counts. LLMs exhibit emergent abilities in reasoning and knowledge use as they grow in size (Wei et al., 2022), and most openly released decoders and their vision-augmented variants are at least several billion parameters. Interestingly, scaling trends are less pronounced for embedding models: while size remains a performance factor, compact encoders often almost rival the performance of much larger ones, indicating diminishing returns with additional parameters (Muennighoff et al., 2022). This suggests it is likely possible to achieve strong visual representation quality with relatively small, well-trained models, especially leveraging late interaction methods (Clavié, 2024).
@@
Recent papers and model releases in the visual retrieval space have claimed performance improvements by scaling the amount of contrastive data and the compute budget (Zhang et al., 2025a; Xu et al., 2025), modifying the attention mask (Chen et al., 2025), increasing image resolutions (Cohere, 2024) or by introducing more diverse tasks and data sources (Jiang et al., 2025). In this work, we attempt to centralize these efforts and systematically disentangle the impact of core design decisions in visual retriever training. Through controlled experiments—ranging from language model pretraining to multi-stage, domain-specific finetuning, we aim to answer a central question:
@@
Contribution 2: ModernVBERT.  Building on these insights, we release ModernVBERT, a small 250M multimodal encoder that aligns a pretrained language encoder with a vision encoder through Masked Language Modeling (MLM) objective, and ColModernVBERT a variant finetuned for document retrieval. Despite its modest size and limited training budget, ColModernVBERT matches models 10x larger on standard visual document retrieval benchmarks, demonstrating the interest of designing a retrieval focused model from the ground up. We publicly release the model, intermediate checkpoints, and the training code at https://huggingface.co/ModernVBERT.
@@ 2 Methodology
Our analysis aims at quantifying the impact of design decisions made when training visual retrievers. In opposition to previous work, we begin our analysis as early as language model modality alignment and iteratively study design choices by modifying design choices independently to reduce confounding factors as much as possible (Allen-Zhu & Li, 2025).
Controlled Experimental Setup. A central point of interest is the impact of causal and bidirectional attention masks. While recently studied for textual representation applications (Gisserot-Boukhlef et al., 2025; Weller et al., 2025), we extend the experiment to the vision modality. We use checkpoints released by Gisserot-Boukhlef et al. (2025) which consist in a series of identical 210M parameter transformer models based on the Llama architecture (Touvron et al., 2023) trained on 100B tokens that differ only in their attention masking strategy during language model training but that are perfectly identical in terms of training data seen, model size and architecture, learning rate scheduling, etc… The checkpoints we use are enc a bidirectional encoder trained with Masked Language Modeling (MLM), dec, a causal decoder trained with next token prediction, and dec-enc a causal decoder annealed over the end of its textual training by removing the causal mask and switching the training objective to MLM. For the vision tower, we employ siglip2-base-16b-512111https://huggingface.co/google/siglip2-base-patch16-512 (Tschannen et al., 2025), a 86M parameter vision transformer contrastively trained on billions of text-image pairs. All ablations thus stem from iso-data controlled setups, and as further described, are further trained on the same data sequence, with the same batch sizes, optimizers, schedulers and on the same hardware.
@@
Model Architecture. Our analysis are not centered around model architectures and to draw broadly applicable insights, we design vision-language models following current standard training practices. In line with most recent work, we employ the early fusion architecture (Alayrac et al., 2022) illustrated in Figure 2, in which visual patch embeddings produced by the vision encoder are projected into the language model input embedding space and concatenated with text token embeddings to encourage joint processing  (Li et al., 2022; Alayrac et al., 2022; Wang et al., 2024; Yang et al., 2025; Marafioti et al., 2025). As described in subsection 2.1, we generalize the training loss to function both with causal and masked language modeling objectives. To handle dynamic resolutions, we split large images into $\times512$ pixel patches as expected by the SigLIP encoder, and further concatenate a downscaled version of the full image to improve inter-patch consistency and global visual understanding following current standard practices  (Lin et al., 2023; Ye et al., 2023). To compress the information from sequences of large images, we apply pixel shuffling (Shi et al., 2016) with a ratio of $r=4$, following prior work on models of comparable size (Marafioti et al., 2025).
Training Procedure. Our experiments focus on retrieval performance. We employ a standard biphasic training procedure, in which we first run modality alignement to train a pretrained textual language model to understand visual inputs through language modeling objectives  (Liu et al., 2023b) (subsection 2.1), then rely on a second text-image contrastive learning phase to learn efficient image representations  (Radford et al., 2021) (subsection 2.2). We further describe the general setup, and detail specific modifications to the default training procedure in the experiment section.
@@ 2.1 Modality Alignment
We align the vision encoder tower with the language model by training the image embedding projection layer to map visual features into the language model embedding space. The pretrained language model is also finetuned with Low-Rank Adapters (LoRA)(Hu et al., 2021), allowing both image and text models to adapt jointly while reducing the risk of monomodal performance collapse (Alayrac et al., 2022; Liu et al., 2023b; Laurençon et al., 2024c; McKinzie et al., 2024; Marafioti et al., 2025).
Alignment Loss. For decoder‑based models, we train with Causal Language Modeling (CLM) loss on the text tokens, as standardly done in VLM modality alignement:
@@
Modality Alignment Corpus. Models are modality aligned on a large corpus in large parts derived from The Cauldron 2 (Laurençon et al., 2024c) and Docmatix (Laurençon et al., 2024a). Our objective being to train document focused retrieval models, we use an adjusted training mixture that upsamples images containing text and documents with varying level of complexities. Our final training corpus consists of approximately 2B text tokens, and includes diverse sources such as web pages, books, and scientific papers. Mixture details are given in Appendix A.3.1. We note that controlling the exact data distribution during this phase enables the models we train to specialize early and achieve good document focused downstream performances which many large models struggle with (Liu et al., 2023a).
Parameters. All models are trained using a masking ratio of 0.5 and user-prompt masking to avoid overfitting on chat-template format (Huerta-Enochian & Ko, 2024; Shi et al., 2024; Allal et al., 2025). We employ WSD scheduler (Hu et al., 2024b) with the first 5% of the training as warmup, the last 20% as decay and a maximum learning rate of 1e-4. The ablation models are aligned on 3.5B tokens. We provide additional details on the training setup in Appendix A.1.
@@ 2.2 Contrastive Post-Training
Once the language model has learned to process image tokens jointly with text tokens, we specialize models through a contrastive post-training stage designed to enhance the semantic representation of the output embeddings produced by the model (Reimers & Gurevych, 2019).
Post-training Pairs. The post-training dataset used as starting point in our ablations comprises 118M document-query pairs from the ColPali corpus Faysse et al. (2025) as well as another 118M of natural image-description pairs from the MSCOCO train set (Lin et al., 2015).
Contrastive Loss. We employ the InfoNCE loss (van den Oord et al., 2019), defined as
@@
where $\mathbf{d^{+}}$ denotes the positive target for the query $\mathbf{q}$, $\mathcal{N}_{\mathbf{q}}=\mathcal{N}_{\mathbf{q}}^{\text{in}}\cup\mathcal{N}_{\mathbf{q}}^{\text{hard}}$ the set of negative targets (in-batch and hard negatives when mentioned), and $\Phi(\mathbf{q},\mathbf{d})$ a similarity function between the token(s) of the query and the documents. [^2] We use the last (EOS) token for causal models, and mean pool all sequence tokens for bidirectional encoders for single-vector models. Alternatively, we use all document and query tokens without pooling for late interaction matching (Faysse et al., 2025). Details in Appendix A.2. For general-domain post-training we compute the loss symmetrically (Radford et al., 2021).
Batches Curation. In contrastive learning, batch diversity critically impacts retrieval entropy. Overly heterogeneous batches lead to trivial retrievals, while curated batches yield richer training signals. We employ task-aware batching (Li et al., 2023), grouping documents by source to ensure a homogeneous batch composition.
@@ 2.3 Ablation Evaluation Setup
The contrastively trained models are evaluated on retrieval and zero-shot classification tasks across multiple domains. Although the main focus remains document retrieval capabilities, evaluated by aggregating scores from the ViDoRe and ViDoRe v2333We report only the English splits of ViDoRe v2, as our base models are trained on English data only. (Macé et al., 2025) benchmarks (nDCG@5), we also assess more generalist image retrieval capabilities by selecting tasks from MIEB (Xiao et al., 2025a). For natural image retrieval, we aggregate MSCOCO retrieval  (Lin et al., 2015) and Flickr30k retrieval (nDCG@10) (Plummer et al., 2016) test sets. Finally, following practices in (Muennighoff et al., 2022), we assess both zero-shot and fine-tuning abilities of our models on general classification tasks. Specifically, we measure classification accuracy by finetuning a logistic regression head on top of our model’s embedding on Stanford Cars (Krause et al., 2013) and Food101 (Bossard et al., 2014), and we evaluate zero-shot performance on FER2013 (Khaireddin & Chen, 2021) and EuroSAT (Helber et al., 2019) and aggregate the results.
@@ 3.1 Modality Alignment Design
Language modeling Modality Alignment improves document understanding. According to benchmarks such as MIEB (Xiao et al., 2025a), dual encoder models explicitly trained on contrastive image-text tasks outperform repurposed VLMs in natural image classification tasks. To assess this, we train an encoder and a decoder vision-language model using the methodology described in section 2 on a mix of natural image and document data (alignment and contrastive training). Then, we compare them with the (unfinetuned) standalone model (SigLIP) used as the vision tower, which has been contrastively trained from scratch on billions of text-image pairs. As shown in Figure 3, the two early fusion VLMs variants severely underperform the dual encoder on natural image tasks. In contrast, they achieve significant gains in document retrieval tasks (+10.9 nDCG@5 on ViDoRe and ViDoRe v2 datasets). This confirms high-level representation tasks do not benefit from the granular interactions between image and text tokens learned during the VLM modality alignment phase, but large-scale contrastive training remains superior. However, pairing a vision model with an LM enables token level interactions that create richer document understanding, aiding specific tasks even with less contrastive post-training.
Scaling the modality alignment phase for better token representations. Prior work shows that scaling the modality alignment phase of VLMs improves their generative abilities (Beyer et al., 2024; McKinzie et al., 2024; Wang et al., 2024). We test whether similar gains hold in retrieval by contrastively finetuning enc checkpoints during MLM modality alignment. Figure 4 illustrates the results of post-trained checkpoints on diverse tasks. Although document retrieval improves consistently with more modality alignment data – largely surpassing the vision tower evaluated in isolation and showing clear scaling benefits – natural image tasks plateau past 1B tokens, far from the standalone dual encoder baseline. This shows that document and natural image retrieval leverage different mechanisms and should not be optimized the same way. Document Retrieval benefits from learning fine-grained interactions between image and text tokens through the language model, while the LM has limited utility for high level natural image tasks.
@@
Bidirectional attention fully unlocks Late Interaction. Inspired by the effectiveness of bidirectional attention in text-only retrieval (Gisserot-Boukhlef et al., 2025; Weller et al., 2025) [^4] Chen et al. (2025) investigate post-hoc removal of the attention mask during visual retrieval finetuning., we investigate if it surpasses causal attention in visual document retrieval, particularly when using the multi-vector late interaction matching common in SOTA visual retrievers (Khattab & Zaharia, 2020; Faysse et al., 2025). Figure 5 reports single vector and late interaction results on the ViDoRe benchmark for various model variants. On top of the standard enc (MLM) and dec (CLM) models, we evaluate the dec-enc and the dec models modality aligned with MLM objectives to determine whether bidirectional attention capabilities can be obtained in later stages of training.
@@
Removing the causal attention mask during training does not suffice to recover the enc late interaction performance at these data regimes. This indicates converting trained decoders as late interaction retrievers is highly non trivial, and confirms the insights from Weller et al. (2025); when possible, training encoder models from scratch remain better for retrieval tasks.

### v2

@@ Abstract
Retrieving specific information from a large corpus of documents is a prevalent industrial use case of modern AI, notably due to the popularity of Retrieval-Augmented Generation (RAG) systems. Although neural document retrieval models have historically operated exclusively in the text space, Visual Document Retrieval (VDR) models – large vision–language decoders repurposed as embedding models which directly work with page screenshots as inputs – are increasingly popular due to the performance and indexing latency gains they offer. In this work, we show that, while cost-efficient, this approach of repurposing generative models bottlenecks retrieval performance. Through controlled experiments, we revisit the entire training pipeline, and establish a principled recipe for improving visual document retrieval models. We notably measure the impact of attention masking, image resolution, modality alignment data regimes, and late interaction centered contrastive objectives which emerge as central performance factors. Building on these insights, we release ModernVBERT, a compact 250M-parameter vision–language encoder that outperforms recent models up to 10 times larger when fine-tuned on document retrieval tasks, enabling efficient inference on cheap CPU hardware and greatly reducing latency and costs while maintaining strong performance. Models, code and data are available at https://huggingface.co/ModernVBERT.
@@ 1 Introduction
The ability to quickly locate specific information in vast document collections is a core building block of digital systems today, supporting use cases that range from web search and virtual assistants to enterprise knowledge management. Neural information retrieval (IR) models, and in particular dense retrievers, have become the de facto backbone of modern search systems thanks to their strong semantic matching capabilities and good scalability properties (reimers_sentence-bert_2019; karpukhin_dense_2020; wang_text_2022).
This trend is amplified by the widespread adoption of Retrieval-Augmented Generation (RAG) (lewis_retrieval-augmented_2020), where a retriever is used to select a small set of relevant documents that condition a downstream generator. In such systems, the first-stage retrieval module is a well-known bottleneck: its recall directly upper-bounds the quality of the generated answers, while its latency and indexing costs partially drive the overall system efficiency (lin-byrne-2022-retrieval). As a result, improving document retrieval, especially for long, complex files such as PDFs, scientific articles, and reports, is a key lever for making industrial RAG deployments more accurate and cost-effective. Visual Document Retrieval. Historically, document retrieval in these settings has operated purely in the text space. To index PDFs or scans, practitioners first run heavy preprocessing pipelines that include Optical Character Recognition (OCR), layout analysis, and heuristic passage segmentation, before embedding the resulting text spans with a neural encoder. This approach suffers from several limitations: OCR and layout parsing can be brittle and slow, complex visual elements such as tables, figures, and typography are often poorly captured, and any error or bias introduced during preprocessing is propagated to the retriever.
Visual Document Retrieval (VDR) has emerged as a compelling alternative to such text-based systems. Rather than indexing pre-extracted textual content, VDR models directly operate on page screenshots: given a user query, they retrieve relevant document pages by matching the query against image-based representations of the pages (faysse2025colpaliefficientdocumentretrieval). By bypassing OCR and layout parsing, VDR yields simpler end-to-end pipelines, significantly reduces indexing latency, and better exploits visual cues such as layout, figures, and fonts, while achieving strong performance on visually rich benchmarks like ViDoRe.
Limits of Generative VLM Repurposing. Most current VDR systems are obtained by repurposing large generative vision–language decoders  (alayrac2022flamingovisuallanguagemodel) as retrieval encoders via post-hoc contrastive fine-tuning (ma2024unifyingmultimodalretrievaldocument; faysse2025colpaliefficientdocumentretrieval; jiang2025vlm2vectrainingvisionlanguagemodels). While cost-efficient, this design choice bottlenecks retrieval performance and efficiency: model sizes, attention patterns, image resolutions, and training objectives are designed for generative use cases rather than optimized for retrieval which has been shown in text models to be suboptimal (lee2025nvembedimprovedtechniquestraining; gisserotboukhlef2025pretrainencodersmaskedlanguage). Furthermore, scaling trends (wei2022emergentabilitieslargelanguage) are less pronounced for embedding models; while correlated with model size, strong retrieval performance remains attainable with small models (clavié2024bettermonolingualjapaneseretrievers).
@@
Recent papers and model releases in the visual retrieval space have claimed performance improvements by scaling the amount of contrastive data and the compute budget (zhang2025gmeimprovinguniversalmultimodal; xu2025llamanemoretrievercolembedtopperforming), modifying the attention mask (chen2025mocamodalityawarecontinualpretraining), increasing image resolutions (cohere_introducing_2024) or by introducing more diverse tasks and data sources (jiang2025vlm2vectrainingvisionlanguagemodels).
In this work, we attempt to centralize these efforts and systematically disentangle the impact of core design decisions in visual retriever training. Through controlled experiments—ranging from language model pretraining to multi-stage, domain-specific fine-tuning, we aim to answer a central question:
@@
Contribution 2: ModernVBERT.  Building on these insights, we release ModernVBERT, a small 250M multimodal encoder that aligns a pretrained language encoder with a vision encoder through Masked Language Modeling (MLM) objective, and ColModernVBERT a variant fine-tuned for document retrieval. Despite its modest size and limited training budget, ColModernVBERT matches models 10x larger on standard visual document retrieval benchmarks, demonstrating the interest of designing a retrieval focused model from the ground up. We release the model, intermediate checkpoints, and the training code at https://huggingface.co/ModernVBERT.
@@ 2 Methodology
Our analysis aims at quantifying the impact of design decisions made when training visual retrievers. In opposition to previous work, we begin our analysis as early as language model modality alignment and iteratively study design choices by modifying design choices independently to reduce confounding factors as much as possible (allenzhu2025physicslanguagemodels1).
Controlled Experimental Setup. A central point of interest is the impact of causal and bidirectional attention masks. While recently studied for textual representation applications (gisserotboukhlef2025pretrainencodersmaskedlanguage; weller2025seqvsseqopen), we extend the experiment to the vision modality. We use checkpoints released by gisserotboukhlef2025pretrainencodersmaskedlanguage which consist in a series of identical 210M parameter transformer models based on the Llama architecture (touvron_llama_2023) trained on 100B tokens that differ only in their attention masking strategy during language model training but that are perfectly identical in terms of training data seen, model size and architecture, learning rate scheduling, etc… The checkpoints we use are enc a bidirectional encoder trained with Masked Language Modeling (MLM), dec, a causal decoder trained with next token prediction, and dec-enc a causal decoder annealed over the end of its textual training by removing the causal mask and switching the training objective to MLM. For the vision tower, we employ the vision component of siglip2-base-16b-512 (tschannen2025siglip2multilingualvisionlanguage), a 86M parameter vision transformer contrastively trained on billions of text-image pairs. All ablations thus stem from iso-data controlled setups, and as further described, are further trained on the same data sequence, with the same batch sizes, optimizers, schedulers and on the same hardware.
@@
Model Architecture. Our analysis are not centered around model architectures and to draw broadly applicable insights, we design vision-language models following current standard training practices. In line with most recent work, we employ the early fusion architecture (alayrac2022flamingovisuallanguagemodel) illustrated in Figure 2, in which visual patch embeddings produced by the vision encoder are projected into the language model input embedding space and concatenated with text token embeddings to encourage joint processing  (li2022blipbootstrappinglanguageimagepretraining; alayrac2022flamingovisuallanguagemodel; wang2024qwen2vlenhancingvisionlanguagemodels; yang2025qwen251mtechnicalreport; marafioti2025smolvlmredefiningsmallefficient). As described in subsection 2.1, we generalize the training loss to function both with causal and masked language modeling objectives. To handle dynamic resolutions, we split large images into $\times512$ pixel tiles as expected by the SigLIP encoder [^1] Images are downscaled (or upscaled) so that the lengths and widths reach a multiple of 512 pixels to preserve the aspect ratio, padding is used on the smaller side when necessary (i.e. a 1024x1000 px image would be padded to 1024x1024 px).. Following current standard practices, we further process a downscaled version of the full image to improve inter-tile consistency and global visual understanding (lin2023sphinxjointmixingweights; ye2023ureaderuniversalocrfreevisuallysituated). The vision tower produces 1024 pixel patch representations for each tile [^2] The SigLIP tower takes 512x512 px images and process them by 16x16 px patches (dosovitskiy_image_2020). This results in $(512/16)^{2}=1024$ patches., which we compress to 64 tokens through pixel shuffling (shi2016realtimesingleimagevideo) with a compression ratio $r=4$, following prior work on models of comparable size (marafioti2025smolvlmredefiningsmallefficient). We highlight the impact of image resolution and this parameter on the number of visual tokens in Appendix C.6.1.
Training Procedure. Our experiments focus on retrieval performance. We employ a standard biphasic training procedure, in which we first run modality alignment to train a pretrained textual language model to understand visual inputs through language modeling objectives  (liu2023visualinstructiontuning) (subsection 2.1), then rely on a second text-image contrastive learning phase to learn efficient image representations  (radford2021learningtransferablevisualmodels) (subsection 2.2). We further describe the general setup, and detail specific modifications to the default training procedure in the experiment section.
@@ 2.1 Modality Alignment
We align the vision encoder tower with the language model by training the image embedding projection layer to map visual features into the language model embedding space. The pretrained language model is also fine-tuned with Low-Rank Adapters (LoRA) (hu2021loralowrankadaptationlarge), allowing both image and text models to adapt jointly while reducing the risk of monomodal performance collapse (alayrac2022flamingovisuallanguagemodel; liu2023visualinstructiontuning; laurençon2024mattersbuildingvisionlanguagemodels; mckinzie2024mm1methodsanalysis; marafioti2025smolvlmredefiningsmallefficient).
Alignment Loss. For decoder‑based models, we train with Causal Language Modeling (CLM) loss on the text tokens, as standardly done in VLM modality alignment:
@@
Modality Alignment Corpus. Models are modality aligned on a large corpus in large parts derived from The Cauldron 2 (laurençon2024mattersbuildingvisionlanguagemodels) and Docmatix (laurençon2024building). Our objective being to train document focused retrieval models, we use an adjusted training mixture that upsamples images containing text and documents with varying level of complexities. Our final training corpus consists of approximately 2B text tokens, and includes diverse sources such as web pages, books, and scientific papers. Mixture details are given in Appendix A.3.1. We note that controlling the exact data distribution during this phase enables the models we train to specialize early and achieve good document focused downstream performances which many large models struggle with (liu_improved_2023).
Parameters. All models are trained using a masking ratio of 0.5 and user-prompt masking to avoid overfitting on chat-template format (huertaenochian2024instructionfinetuningdoesprompt; shi2024instructiontuninglossinstructions; allal2025smollm2smolgoesbig). We employ WSD scheduler (hu2024minicpmunveilingpotentialsmall) with the first 5% of the training as warmup, the last 20% as decay and a maximum learning rate of 1e-4. The ablation models are aligned on 3.5B tokens. We provide additional details on the training setup in Appendix A.1.
@@ 2.2 Contrastive Post-Training
Once the language model has learned to process image tokens jointly with text tokens, we specialize models through a contrastive post-training stage designed to enhance the semantic representation of the output embeddings produced by the model (reimers_sentence-bert_2019).
Post-training Pairs. The post-training dataset used as starting point in our ablations comprises 118k document-query pairs from the ColPali corpus faysse2025colpaliefficientdocumentretrieval as well as another 118k of natural image-description pairs from the MSCOCO train set (lin2015microsoftcococommonobjects).
Contrastive Loss. We employ the InfoNCE loss (oord2019representationlearningcontrastivepredictive), defined as
@@
where $\mathbf{d^{+}}$ denotes the positive target for the query $\mathbf{q}$, $\mathcal{N}_{\mathbf{q}}=\mathcal{N}_{\mathbf{q}}^{\text{in}}\cup\mathcal{N}_{\mathbf{q}}^{\text{hard}}$ the set of negative targets (in-batch and hard negatives when mentioned), and $\Phi(\mathbf{q},\mathbf{d})$ a similarity function between the token(s) of the query and the documents. [^3] We use the last (EOS) token for causal models, and mean pool all sequence tokens for bidirectional encoders for single-vector models. Alternatively, we use all document and query tokens without pooling for late interaction matching (faysse2025colpaliefficientdocumentretrieval). Details in Appendix A.2. For general-domain post-training we compute the loss symmetrically (radford2021learningtransferablevisualmodels).
Batches Curation. In contrastive learning, batch diversity critically impacts retrieval entropy. Overly heterogeneous batches lead to trivial retrievals, while curated batches yield richer training signals. We employ task-aware batching (li2023generaltextembeddingsmultistage), grouping documents by source to ensure a homogeneous batch composition.
@@ 2.3 Ablation Evaluation Setup
The contrastively trained models are evaluated on retrieval and zero-shot classification tasks across multiple domains. Although the main focus remains document retrieval capabilities, evaluated by aggregating scores from the ViDoRe and ViDoRe v2444We report only the English splits of ViDoRe v2, as our base models are trained on English data only. (macé2025vidorebenchmarkv2raising) benchmarks (nDCG@5), we also assess more generalist image retrieval capabilities by selecting tasks from MIEB (xiao2025miebmassiveimageembedding). For natural image retrieval, we aggregate MSCOCO retrieval  (lin2015microsoftcococommonobjects) and Flickr30k retrieval (nDCG@10) (plummer2016flickr30kentitiescollectingregiontophrase) test sets. Finally, following practices in (muennighoff_mteb_2022), we assess both zero-shot and fine-tuning abilities of our models on general classification tasks. Specifically, we measure classification accuracy by fine-tuning a logistic regression head on top of our model’s embedding on Stanford Cars (6755945) and Food101 (10.1007/978-3-319-10599-4_29), and we evaluate zero-shot performance on FER2013 (khaireddin2021facialemotionrecognitionstate) and EuroSAT (helber2019eurosatnoveldatasetdeep) and aggregate the results.
@@ 3 What Makes a Great Visual Retriever?
Figure 3: Impact of Modality Alignment objective on downstream tasks. Early Fusion of vision and text models boosts document retrieval tasks regardless of the LM objective, but degrades natural image and classification tasks w.r.t. the standalone fine-tuned vision model SigLIP. Reported scores are aggregated MIEB scores (nDCG, Accuracy.)
@@ 3.1 Modality Alignment Design
Language modeling Modality Alignment improves document understanding. According to benchmarks such as MIEB (xiao2025miebmassiveimageembedding), dual encoder models explicitly trained on contrastive image-text tasks outperform repurposed VLMs in natural image classification tasks. To assess this, we train an encoder and a decoder vision-language model using the methodology described in section 2 on a mix of natural image and document data (alignment and contrastive training). We compare them with SigLIP2-FT, the 378M dual vision encoder model whose vision component is used by the vision tower of both VLMs, and with the larger SigLIP2-FT Large (881M parameters). Both SigLIP-FT models are finetuned in the same conditions as the VLMs, and initialized from pre-trained weights from scratch on billions of text-image pairs [^5] We report the performance of the untrained off-the-shelf SigLIP in Appendix C.1. As shown in Figure 3, the two early fusion VLM variants severely underperform the SigLIP2-FT dual encoders on natural image tasks. In contrast, they achieve significant gains on document retrieval tasks (+6.1 nDCG@5 on ViDoRe and ViDoRe v2 datasets w.r.t. base), even edging out SigLIP2-FT Large that contains 3.5x vision parameters more than both VLMs.
This confirms large-scale contrastive training remains best for high-level image representation tasks (natural images), but sequentially combining a vision model with a pretrained language model facilitates document representation tasks, even with significantly less contrastive post-training. As the rest of this paper shows, steering away from the dual encoder architecture further enables improving performance through many avenues other than text to image contrastive training, for which supervised training samples can be hard to obtain.
Scaling the modality alignment phase for better token representations. Prior work shows that scaling the modality alignment phase of VLMs improves their generative abilities (beyer2024paligemmaversatile3bvlm; mckinzie2024mm1methodsanalysis; wang2024qwen2vlenhancingvisionlanguagemodels). We test whether similar gains hold in retrieval by contrastively fine-tuning enc checkpoints during MLM modality alignment. Figure 4 illustrates the results of post-trained checkpoints on diverse tasks. Although document retrieval improves consistently with more modality alignment data – largely surpassing the vision tower evaluated in isolation and showing clear scaling benefits – natural image tasks plateau past 1B tokens, far from the standalone dual encoder baseline. This shows that document and natural image retrieval leverage different mechanisms and should not be optimized the same way. Document Retrieval benefits from learning fine-grained interactions between image and text tokens through the language model, while the LM has limited utility for high level natural image tasks.
@@
Bidirectional attention fully unlocks Late Interaction. Inspired by the effectiveness of bidirectional attention in text-only retrieval (gisserotboukhlef2025pretrainencodersmaskedlanguage; weller2025seqvsseqopen)666chen2025mocamodalityawarecontinualpretraining investigate post-hoc removal of the attention mask during visual retrieval fine-tuning., we investigate if it surpasses causal attention in visual document retrieval, particularly when using the multi-vector late interaction matching common in SOTA visual retrievers (khattab_colbert_2020; faysse2025colpaliefficientdocumentretrieval). Figure 5 reports single vector and late interaction results on the ViDoRe benchmark for various model variants. On top of the standard enc (MLM) and dec (CLM) models, we evaluate the dec-enc and the dec models modality aligned with MLM objectives to determine whether bidirectional attention capabilities can be obtained in later stages of training.
@@
Removing the causal attention mask during training does not suffice to recover the enc late interaction performance at these data regimes. This indicates converting trained decoders as late interaction retrievers is highly non trivial, and confirms the insights from weller2025seqvsseqopen; when possible, training encoder models from scratch remain better for retrieval tasks.